tristan22:

she descends 18 feet in 3 minutes

Eiwoh2:

Welcome to QuestionCove! Could you give us the full question to understand? :)

tristan22:

okay

tristan22:

cami goes scuba diving she descends 18 feet in 3minutes which expressions represent the average rate at which cami position changes in feet per minute?

Hoodmemes:

Can u post the following expressions?

tristan22:

okay

tristan22:

-18/-3 - (18/3) - 18/3 6 -6

Hoodmemes:

-(18/3) = -1 x (18/3) = -1(6) = (blank)